# “I Want It That Way”: Enabling Interactive Decision Support Using Large Language Models and Constraint Programming

### v1

@@ 4.3. Constraint Programming Solver
In response to feedback from the diary study, the MeetMate system was also designed to return diverse time suggestions. Instead of returning only a single-time suggestion with the best weighted score, we return a set of $k$ diverse times. We define diversity as the sum of the pairwise distance between each time in the returned set $\{t_{1},\dots,t_{k}\}$:
@@
where $d(t_{1},t_{2})$ is a distance function between two suggested times. Motivated by the results of the diary study which highlighted the importance of temporal diversity, we define the distance between two candidate times as the logarithm of their difference in minutes plus one. Finding diverse subsets is known to be a computationally demanding problem (Agrawal et al., 2009). To construct diverse time suggestions we employ a simple greedy algorithm that we found works well in practice. We start by filtering the candidate times to only times within $\epsilon$ of the best score. We select $\epsilon$ to be the smallest value such that there are at least $k$ times that meet the criteria. After filtering the times, we greedily construct the set by first selecting the earliest high-scoring time and iteratively adding a new time to the set that maximizes the diversity of the current set.
To help provide transparency over the suggested meeting times, we return each meeting time with a one-sentence summary of its pros and cons. To generate this summary, we give an LLM the list of scheduling preferences met and unmet by each suggested time and prompt it to generate a one-sentence summary. For instance here is a sample explanation for a time that met two user preferences but failed to include one attendee: This time is before 11am on Tuesday, but Anton cannot attend. While LLMs are known to hallucinate and generate text not supported by the inputs (Bubeck et al., 2023), we found that summarizing small lists of preferences rarely led to hallucinations.
@@ 4.4. System Limitations
This system was designed as a prototype to test the viability of our hybrid LLM and constraint programming framework, not as a full calendar scheduling system. As such, we include only the minimum viable amount of functionality to test the interactive component between the user and the system and exclude any additional functionality that may be present in other calendar tools. When scheduling a meeting, our system requires the meeting characteristics to be input in a form as opposed to a naturalistic chat. These characteristics are fixed for the remainder of the interaction (i.e., the user cannot dynamically change the attendees or length of the meeting during chat). The system only supports meetings that begin at increments of 15 minutes (i.e., on the hour, and fifteen-minute increments thereafter). The system is also limited in what it can communicate to users. Specifically, it can communicate suggested meeting times, explanations of those times, and some limited follow-up questions. The system can not interface with external tools or APIs (e.g., sending e-mails to participants, querying weather data, or accessing conference room availability).
@@ 5.1. Datasets
To quantitatively evaluate the system components we create a novel evaluation benchmark that leverages both a synthetic calendar universe and real scheduling preferences captured during the diary study. The synthetic calendar universe is used to create sample meeting scheduling scenarios without the use of users’ personally identifiable private calendar data. To generate the calendar universe we use GPT4 (OpenAI, 2023) to create a synthetic organization with 32 employees spread over 4 distinct teams. We also use the LLM to generate a suite of existing meetings between the employees to populate their calendars. The number of meetings for each employee were commensurate with their position with managers having a higher meeting load, and calibrated to be in line with the self-reported number of attended meetings in the diary study. We also generate a dataset of 75 new meeting instances that represent meetings that need to be scheduled.
@@
To generate a dataset of scheduling preferences we process the results of the diary study. We first extracted all scheduling preferences and constraints elicited during the diary study and removed any inadvertantly shared personally identifiable information (PII). To help generalize the preferences elicited during the diary study we also remove any specific times, days of the week, or meeting specific information (i.e. attendees, project title) and replace them with special placeholders. These placeholders are later in-filled by either sampling the inputs from a uniform distribution over potential inputs (i.e., randomnly sample one of the five different workdays) or entering the relevant meeting characteristic (i.e., inputting the name of the organizer for the meeting instance). For every original preference from the diary study, we generate 3 new preferences by in-filling the placeholder values for new meeting scheduling instances. Figure 6 shows a sample processing pipeline for a final scheduling preference.
Figure 6. Processing pipeline for preferences from the diary study.
@@
We categorize all the processed preferences from the diary study into two categories: requests that can be handled given the limitations of the synthetic calendar universe and those that cannot. For instance, since the synthetic universe does not have any facility information, any preference related to the availability of conference rooms is marked as not handled. These categories form the basis of a binary classification dataset, which we call the Safeguard Dataset, through which we evaluate the information checker component of the system. We also take the subset of the dataset that can be handled by the current system and use it to form a dataset, called the Code Generation Dataset, for the coder component.
@@ 5.2. Information Checker
We use the Safeguard dataset to benchmark the performance of current LLMs for the Information Checker component of the MeetMate system. We evaluate two different LLMs: GPT-3 (text-davinci-003) (Brown et al., 2020) and GPT-4 (OpenAI, 2023). We also evaluate two different phrasing strategies. During initial experiments we noticed that the LLMs struggled to extract correct scheduling preferences when given long sentences with rationale. We add a LLM-powered rephraser component that, given the initial scheduling preference, is asked to rephrase it as a succinct scheduling constraint. We note that this is not necessary in the full implementation of the MeetMate system as the constraint manager component rephrases the chat message into scheduling constraints. We report the classification accuracy of the models and phrasing strategies on the reference dataset, defined as the fraction of the instances the model correctly predicts can and cannot be handled by the current system.
Table 1 (Safeguard Accuracy column) summarizes the accuracy for the two different models and prompting strategies. Both models have around 80% accuracy for the information checking task when using the rephraser. The rephraser does indeed lead to a small increase in performance for both models, highlighting the importance of translating complex user queries into clear requests before generating code. Interestingly, GPT-3 slightly outperforms GPT-4. We hypothesize that this discrepancy is due to GPT-4 over-estimating its ability to handle constraint requests.
@@ 5.3. Code Generation
To evaluate the coder component of the MeetMate system we compare the LLM generated code to ground-truth correct implementations of each function generated by human software developers. The correct implementations follow a similar processing strategy as the original dataset. An implementation is generated only for each processed preference with placeholders from the diary study, and the placeholders are in-filled for the specific meeting context or sampled values. We evaluate three different metrics for the code generation component:
@@ 6. User Study
To evaluate the feasibility of MeetMate  we conducted a user study with 10 participants who used the MeetMate user interface to schedule meetings as a member of a fictitious organization. In this user study, we evalute the MeetMate system not as a finished product but as a probe (Boehner et al., 2007) to observe the interactive preference elicitation process in real-world scheduling scenarios and to uncover design implications and unforeseen challenges for further improving the system design. Our study tasks are, therefore, intentionally designed to push the limits of the current system design.
@@ 6.1. MeetMate UI
We prototyped a UI for MeetMate to allow users to interact with the underlying system. The main goal of the prototype is to provide a minimal set of capabilities that lets us exercise the underlying system components rather than to provide a full set of functionalities of a typical calendar. The UI consists of a calendar component on the left (Figure 7A) and a chatbot component on the right (Figure 7B). The calendar component resembles a typical calendar client and displays the current user’s weekly schedule. The chatbot component allows for naturalistic conversations with the system with a chat history and a message box.
The prototype UI was designed to support only allowable interactions with the underlying components. As such, the user started the interaction by first specifying a list of attendees and the duration of the desired meeting (Figure 7C), and the system responded with an initial suggestion with a short explanation (Figure 7E). All meetings were set to be scheduled in a default time horizon of one week. If the suggested time is satisfactory, the user can schedule the meeting by clicking on the “Schedule” button. If the suggested time is unsatisfactory, the user can freely interact with the system through the message box (Figure 7E). Each time a user sends a message, the system responds with a suggested time and/or a message, in which case the user either accepts the suggestion or iterates by interacting with the system. The system could be configured to return multiple suggestions at a time (Figure 7F).
@@ 6.2. Study Design
Each study session consisted of four scheduling tasks crafted to challenge the system’s capabilities. The tasks and the calendars that participants interacted with were set up either to make it difficult for the system to satisfy all preferences or to allow users to engage with the system in unanticipated ways that would potentially break the system. The first task involved a 40-minute meeting with three other attendees with preferences for it being after 11am, a 30-minute break before the meeting, and allowing for a meeting over lunch, which were preferences that MeetMate  was designed to support. The task was designed to also test the Information Checker by adding a preference for a sunny day, which the system was not designed to support. The second task involved scheduling an urgent, next-day, 60-minute meeting with two required attendees and two optional attendees, including the organizer. The third task involved scheduling a 60-minute meeting with three other attendees, and the participant was allowed to bring in any meeting preferences or constraints they wanted to exercise to evaluate the system. The fourth task was identical to the third task, except that the first three tasks returned one suggestion at a time and the fourth task returned three suggestions at a time. The third task was optional if there was not enough time to complete all four tasks within one hour. Each task was considered done when the participant found a satisfactory meeting time or when they expressed that they could not complete the task. With the exception of the first task, all tasks were designed to be challenging to complete with the current system to help expose new challenges within the interactive decision support framework. Participants were not taught about the limitations of the systems and what scenarios were not supported (e.g., calendar question and answering, dynamically changing meeting time). This was to gauge the types of features participants expected in a dynamic setting and to evaluate how well participants could recover from shortcomings in the system.
@@
Throughout the study, participants were asked to verbalize their thoughts (i.e., think-aloud protocol). During the tasks, the facilitators asked follow-up questions probing the participants’ comments or behaviors. After each task, we asked participants about their ability to find a suitable meeting time, the ease of expressing preferences, the ease of using the system to schedule a meeting, and their understanding of the limits of the system’s abilities. We also asked them for overall feedback about their scheduling experience, a comparison between their current scheduling practice and using MeetMate, and any feedback about the number of suggestions provided.
@@ 6.4. Findings
Overall, users found the interactive decision support system easy to use and responsive to their preferences. Specifically, users highlighted the ease at which they could communicate their preferences to the system: “The natural language input was easy to use. Right. I don’t have to learn syntax or much of anything (P6)”, “I would schedule this way. I’m real partial to chatting, so I’m very used to it (P4)”. Users also remarked how flexible the agent was in responding to their preferences: “When I picked only the first names it was able to understand and that’s a great thing (P9)”, “It totally understood what I meant, so the suggestions it gave me actually works. (P10)”, “I liked that it was able to meet preferences (P5)”. Multiple users expressed wanting to use the current system further (P1, P4, P8).
Our findings also underscored the importance of an interactive and contextualized system for decision support. Participants noted that some scheduling preferences were only revealed within the iterative process of scheduling the meeting. For instance, throughout their interview P7 asserted that they prefer scheduling meetings at the first available time for all required attendees. However, at the end of the interview, they revealed that they “subconsciously prioritize” to avoid booking out their entire calendar, e.g. “I don’t really want to do; they’re only available at 5PM. Yeah, I got in at seven. I’m not doing a 5:30PM meeting (P7).” We saw that this fluid and subliminal nature of preferences was a large reason why several participants (P1, P2, P3, P6, P7, P8, P10) preferred receiving multiple suggestions. P6 elaborated on the benefit of having options to help refine preferences: “Maybe none of them are perfect, but maybe they’re close. Is there a compromise?” Participants also reevaluated preferences based on contextual information in the scheduling tasks. For instance, P10 expressed a preference for a 15 minute break before the meeting if they’ve been in back-to-back meetings for at least 2 hours. When asked if this is how they typically handle breaks, they described scenarios where they would relax this constraint, including “if it’s just 2 hours and then after that I’m completely free” or “if one of those 2 hour meetings is one where I’m just listening in.”

### v3

@@ 4.3. Constraint Programming Solver
In response to feedback from the diary study, the MeetMate system was also designed to return diverse time suggestions. Instead of returning only a single-time suggestion with the best weighted score, we return a set of $k$ diverse times. Note that simply returning the top $k$ times in terms of score, without considering diversity, can result in similar time suggestions as similar times often have similar properties and scores (e.g., suggesting 9-9:30, 9:30-10, and 10-10:30 if they are the first three available time slots). We define diversity as the sum of the pairwise distance between each time in the returned set $\{t_{1},\dots,t_{k}\}$:
@@
where $d(t_{1},t_{2})$ is a distance function between two suggested times. Motivated by the results of the diary study which highlighted the importance of temporal diversity, we define the distance between two candidate times as the logarithm of their difference in minutes plus one. Finding diverse subsets is known to be a computationally demanding problem (Agrawal et al., 2009). To construct diverse time suggestions we employ a simple greedy algorithm that we found works well in practice. We start by filtering the candidate times to only times within $\epsilon$ of the best score (i.e., to ensure high-quality time suggestions). We select $\epsilon$ to be the smallest value such that there are at least $k$ times that meet the criteria. In practice, since our scoring system gives integral scores (i.e., because $w_{i}$ is integer, and $f_{i}(t)$ is binary), there are often many time slots that receive the same score. Intuitively, in settings with fewer scheduling constraints (e.g., three attendees, must be in the afternoon) many potential meeting times often meet the same or similar subsets of constraints (e.g., times on tuesday or thursday afternoon). Consequently, although our setting of $\epsilon$ only guarantees at least $k$ times are included in the filtered set of candidate times, in most instances there are many more times than $k$ included. This leads to a diverse pool of candidate times that are approximately optimal with respect to the given constraints. After filtering the times, we greedily construct the set by first selecting the earliest high-scoring time and iteratively adding a new time to the set that maximizes the diversity of the current set.
To help provide transparency over the suggested meeting times, we return each meeting time with a one-sentence summary of its pros and cons. To generate this summary, we give an LLM the list of scheduling preferences met and unmet by each suggested time and prompt it to generate a one-sentence summary. For instance here is a sample explanation for a time that met two user preferences but failed to include one attendee: This time is before 11am on Tuesday, but Anton cannot attend. While LLMs are known to hallucinate and generate text not supported by the inputs (Bubeck et al., 2023), we found that summarizing small lists of preferences were consistently faithful, and saw no instances of hallucination for this component in the user study.
@@ 4.4. System Limitations
This system was designed as a technology probe (Boehner et al., 2012; Hutchinson et al., 2003) to test the viability of our hybrid LLM and constraint programming framework, not as a full calendar scheduling system. As such, we include only the minimum viable amount of functionality to test the interactive component between the user and the system and exclude any additional functionality that may be present in other calendar tools. When scheduling a meeting, our system requires the meeting characteristics to be input in a form as opposed to a naturalistic chat. Future research is needed to determine whether these meeting characteristics (or a subset thereof) can be elicited robustly from chat and if that provides a better user experience than this GUI-based approach. These meeting characteristics are also fixed for the remainder of the interaction (i.e., the user cannot dynamically change the attendees or length of the meeting during chat). The system only supports meetings that begin at increments of 15 minutes (i.e., on the hour, and fifteen-minute increments thereafter). The system is also limited in what it can communicate to users. Specifically, it can communicate suggested meeting times, explanations of those times, and some limited follow-up questions. The system can not interface with external tools or APIs (e.g., sending e-mails to participants, querying weather data, or accessing conference room availability). We also note that the diversity selection criteria for returning multiple candidate times is fixed and does not allow a user to change what diverse options are meaningful to help their scheduling or specify what constraints they are willing to sacrifice for diverse options (e.g., if a user is willing to compromise on constraints to have meeting time options on different days of the week).
@@ 5.1. Datasets
To quantitatively evaluate the system components we create a novel evaluation benchmark that leverages both a synthetic calendar universe and real scheduling preferences captured during the diary study. The synthetic calendar universe is used to create sample meeting scheduling scenarios without the use of users’ personally identifiable private calendar data. To generate the calendar universe we use GPT4 (Achiam et al., 2023) to create a synthetic organization with 32 employees spread over 4 distinct teams. We also use the LLM to generate a suite of existing meetings between the employees to populate their calendars. The number of meetings for each employee were commensurate with their position with managers having a higher meeting load, and calibrated to be in line with the self-reported number of attended meetings in the diary study. We also generate a dataset of 75 new meeting instances that represent meetings that need to be scheduled.
@@
To generate a dataset of scheduling preferences we process the results of the diary study. We first extracted all scheduling preferences and constraints elicited during the diary study and removed any inadvertantly shared personally identifiable information. To help generalize the preferences elicited during the diary study we also remove any specific times, days of the week, or meeting specific information (i.e., attendees, project title) and replace them with special placeholders. These placeholders are later in-filled by either sampling the inputs from a uniform distribution over potential inputs (e.g., randomnly sample one of the five different workdays) or entering the relevant meeting characteristic (e.g., inputting the name of the organizer for the meeting instance). For every original preference from the diary study, we generate 3 new preferences by in-filling the placeholder values for new meeting scheduling instances. Figure 8 shows a sample processing pipeline for a final scheduling preference.
Figure 8. Processing pipeline for preferences from the diary study. A preference from the diary study (original preference) is pre-processed to replace any specific times, days of the week, or meeting specific information with a special placeholder (e.g., “[Weekday]” in diagram). To generate a set of preferences for the dataset, the placeholder in the processed preference is in-filled with either relevant meeting details from the synthetic scheduling instance, or in-filled uniformly at random from a set of allowable settings (i.e., Monday through Friday for the weekday tag).
[…]
Three components that feed into each other, from Original Preference to Processed Preference to In-Filled Preferences. An example of an Original Preference is ‘My team has a no-meeting policy on Friday.’ This becomes the Processed Preference ‘My team has a no-meeting policy on [WEEKDAY],’ which turns into the In-Filled Preferences ‘My team has a no-meeting policy on [Thursday, Monday, or Friday],’ where Thursday, Monday, and Friday are possible options for [WEEKDAY].
@@
We categorize all the processed preferences from the diary study into two categories: requests that can be handled given the limitations of the synthetic calendar universe and those that cannot. For instance, since the synthetic universe does not have any facility information, any preference related to the availability of conference rooms is marked as not handled. These categories form the basis of a binary classification dataset, which we call the Safeguard Dataset, through which we evaluate the information checker component of the system. We also take the subset of the dataset that can be handled by the current system (i.e., could be implemented in python with available data sources) and use it to form a dataset, called the Code Generation Dataset, for the coder component.
To summarize the overall dataset generation process we first process the diary study results to create a set of natural language scheduling preferences and use it to create two datasets. The Safeguard Dataset is a binary classificaiton dataset that contains natural language constraints and whether or not the current system has sufficient data sources to handle it. The Code Generation Dataset contains a set of natural language contraints which can be implemented in python with the available data structures.
@@ 5.2. Information Checker
We use the Safeguard dataset to benchmark the performance of current LLMs for the Information Checker component of the MeetMate system. We evaluate two different LLMs: GPT-3 (text-davinci-003) (Brown et al., 2020) and GPT-4 (Achiam et al., 2023). We also evaluate two different phrasing strategies. During initial experiments we noticed that the LLMs struggled to extract correct scheduling preferences when given long sentences with rationale. We add a LLM-powered rephraser component that, given the initial scheduling preference, is asked to rephrase it as a succinct scheduling constraint. We note that this is not necessary in the full implementation of the MeetMate system as the constraint manager component rephrases the chat message into scheduling constraints. We report the classification accuracy of the models and phrasing strategies on the reference dataset, defined as the fraction of the instances the model correctly predicts can and cannot be handled by the current system.
Table 1 (Safeguard Accuracy column) summarizes the accuracy for the two different models and prompting strategies. Both models have around 80% accuracy for the information checking task when using the rephraser. The rephraser does indeed lead to a small increase in performance for both models, highlighting the importance of translating complex user queries into clear requests before generating code.
@@ 5.3. Code Generation
To evaluate the coder component of the MeetMate system we compare the LLM generated code for a single preference to implementations of each function generated by human software developers. To ensure the correctness of the implementations, all reference implementations were checked independently by a different software engineer then the one that implemeneted it. The correct implementations follow a similar processing strategy as the original dataset. An implementation is generated only for each processed preference with placeholders from the diary study, and the placeholders are in-filled for the specific meeting context or sampled values. We evaluate three different metrics for the code generation component:
@@ 6. User Study
To evaluate the feasibility of MeetMate, we conducted a user study with 10 participants who used the MeetMate user interface to schedule meetings as a member of a fictitious organization. In this user study, we evalute the MeetMate system not as a finished product but as a “probe” that inspires design ideas (Boehner et al., 2007). The concept of using a probe as a provocational or an inspirational instrument to “open a conversation (Boehner et al., 2012)” and to “find out about the unknown—to hopefully return with useful or interesting data (Hutchinson et al., 2003)” has been well-established in the HCI literature for many decades (Gaver et al., 1999; Graham and Rouncefield, 2008; Boehner et al., 2007; Hutchinson et al., 2003). This method invites user participation in the early stages of the design process and to promote flexible and open-ended use with less emphasis on usability evaluation. In this user study, we utilize the MeetMate  as a “technology probe”—a functional piece of technology artifact presented to the users—with social science, engineering, and design goals as outlined by (Hutchinson et al., 2003). Our goal with the technology probe is to observe the interactive preference elicitation process in real-world scheduling scenarios and to uncover design implications and unforeseen challenges for further improving the system design. Our study protocol and tasks are, therefore, intentionally designed to provoke open-ended reactions rather than task successes and to push the limits of the current system design.
@@ 6.1. MeetMate UI
We designed a UI for MeetMate to allow users to interact with the underlying system. The main goal of the design is to provide a minimal set of capabilities that lets us exercise the underlying system components rather than to provide a full set of functionalities of a typical calendar. The UI consists of a calendar component on the left (Figure 9A) and a chatbot component on the right (Figure 9B). The calendar component resembles a typical calendar client and displays the current user’s weekly schedule. The chatbot component allows for naturalistic conversations with the system with a chat history and a message box. All components of MeetMate (i.e., Constraint Manager, Information Checker, and Coder) used GPT-3 for the user study as it had comparable performance to GPT-4 in the quantitative evaluation outlined in Section 5 and, at the time of the user study, had a lower latency than using GPT-4.
The system UI was designed to support only allowable interactions with the underlying components. As such, the user started the interaction by first specifying a list of attendees and the duration of the desired meeting (Figure 9C), similar to modern calendaring tools like Microsoft Outlook or Google calendar, and the system responded with an initial suggestion with a short explanation (Figure 9E). All meetings were set to be scheduled in a default time horizon of one week. If the suggested time is satisfactory, the user can schedule the meeting by clicking on the “Schedule” button. If the suggested time is unsatisfactory, the user can freely interact with the system through the message box (Figure 9E). Each time a user sends a message, the system responds with a suggested time and/or a message, in which case the user either accepts the suggestion or iterates by interacting with the system. The system could be configured to return multiple suggestions at a time (Figure 9F).
@@
[…]
Two components: (A) Calendar and (B) Chatbot. Chatbot subcomponents can include a (C) Scheduling Form, (D) Initial Time Suggestion, (E) User Input, and (F) Multiple Time Suggestions.
@@ 6.2. Study Design
Each study session consisted of four scheduling tasks crafted to challenge the system’s capabilities. The tasks and the calendars that participants interacted with were set up either to make it difficult for the system to satisfy all preferences or to allow users to engage with the system in unanticipated ways that would potentially break the system. The first task involved a 40-minute meeting with three other attendees with preferences for it being after 11am, a 30-minute break before the meeting, and allowing for a meeting over lunch, which were preferences that MeetMate  was designed to support. The task was designed to also test the Information Checker by adding a preference to meet on a sunny day, which the system was not designed to support. The second task involved scheduling an urgent, next-day, 60-minute meeting with two required attendees and two optional attendees, including the organizer. The third task involved scheduling a 60-minute meeting with three other attendees, and the participant was allowed to bring in any meeting preferences or constraints they wanted to exercise to evaluate the system. The fourth task was identical to the third task, except that the first three tasks returned one suggestion at a time and the fourth task returned three suggestions at a time. The modality of the meetings (e.g., virtual, in person, hybrid) were not explicitly stated in the user study instructions giving users the maximum freedom to determine what scheduling constraints they wanted to consider. The third task was optional if there was not enough time to complete all four tasks within one hour (3 out of 10 participants completed it). Note that in all tasks, users were provided with the duration, time frame, and attendees of the meeting they were instructed to schedule. Each task was considered done when the participant found a satisfactory meeting time or when they expressed that they could not complete the task. With the exception of the first task, all tasks were designed to be challenging to complete with the current system to help expose new challenges within the interactive decision support framework. Participants were not taught about the limitations of the systems and what scenarios were not supported (e.g., calendar question and answering, dynamically changing meeting time). This was to gauge the types of features participants expected in a dynamic setting and to evaluate how well participants could recover from shortcomings in the system.
@@
Throughout the study, participants were asked to verbalize their thoughts (i.e., think-aloud protocol). During the tasks, the facilitators asked follow-up questions probing the participants’ comments or behaviors. After each task, we asked participants about their ability to find a suitable meeting time, the ease of expressing preferences, the ease of using the system to schedule a meeting, and their understanding of the limits of the system’s abilities. We also asked them for overall feedback about their scheduling experience, a comparison between their current scheduling practice and using MeetMate, and any feedback about the number of suggestions provided. These questions were used as a starting point of open-ended discussions about the system design and the tasks. It is important to note that, while our study allowed for capturing of quantifiable data such as task success rate, we do not report such metrics because of the inherent open-ended and conversational nature of our study design that make such metrics invalid.
@@ 6.4. Findings
Overall, users found the interactive decision support system easy to use and responsive to their preferences. Specifically, users highlighted the ease at which they could communicate their preferences to the system: “The natural language input was easy to use. Right. I don’t have to learn syntax or much of anything (P6)”, “I would schedule this way. I’m real partial to chatting, so I’m very used to it (P4)”. Users also remarked how flexible the agent was in responding to their preferences: “When I picked only the first names it was able to understand and that’s a great thing (P9)”, “It totally understood what I meant, so the suggestions it gave me actually works. (P10)”, “I liked that it was able to meet preferences (P5)”. The system also supported evolving preferences over time. For instance, P2 was able to instruct the system to find a time that could include optional meeting participants at the expense of satisfying a preference for having the meeting the next calendar day. Multiple users expressed wanting to use the current system further (P1, P4, P8).
Our findings also underscored the importance of an interactive and contextualized system for decision support. Participants noted that some scheduling preferences were only revealed within the iterative process of scheduling the meeting. For instance, throughout their interview P7 asserted that they prefer scheduling meetings at the first available time for all required attendees. However, at the end of the interview, they revealed that they “subconsciously prioritize” to avoid booking out their entire calendar, for instance, “I don’t really want to do [book this time]; they’re only available at 5PM. Yeah, I got in at 7 [AM]. I’m not doing a 5:30PM meeting (P7).” We saw that this fluid and subliminal nature of preferences was a large reason why several participants (P1, P2, P3, P6, P7, P8, P10) preferred receiving multiple suggestions. P6 elaborated on the benefit of having options to help refine preferences: “Maybe none of them are perfect, but maybe they’re close. Is there a compromise?” Participants also reevaluated preferences based on contextual information in the scheduling tasks. For instance, P10 expressed a preference for a 15 minute break before the meeting if they have been in back-to-back meetings for at least 2 hours. When asked if this is how they typically handle breaks, they described scenarios where they would relax this constraint, including “if it’s just 2 hours and then after that I’m completely free” or “if one of those 2 hour meetings is one where I’m just listening in.”